OpenStudy (anonymous):

Consider the functions f(x) = 2(x − 3) and g(x) = 3x. What is the value of (f − g)(3)?

OpenStudy (cwrw238):

(f-g) = 2x-6 - 3x = -x-6 )f-g)(3) = -3-6 = -9

OpenStudy (cwrw238):

are you ok with that ?